OpenStudy (anonymous):

The slope of a line is –2 and its y-intercept is (0, 3). What is the equation of the line that is parallel to the first line and passes through (2, 2)?

OpenStudy (anonymous):

@yosup if both lines are paralell then thier slope is same so eq of line paralell to it and with oint (2,2) is (y-y1)=m(x-x1) =>(y-2)=-2(x-2) =>y-2=-2x+4 =>2x+y=6 or inn standard form y=-2x+6